In the Cartesian plane, two vertices of a square have coordinates (3 ; 4) and ( -2 ; - 1). One of the other two vertices, has coordibates


1
Expert's answer
2022-02-22T10:45:00-0500

Two opposite vertices of a square are "(3,4)" and "(-2,-1)." Find the coordinates of other two vertices.

Assume "A(-2, -1)" and "C(3, 4)" are two opposite vertices of a square 


"\overrightarrow{AC}=(3-(-2), 4-(-1))"

"|\overrightarrow{AC}|=\sqrt{5^2+5^2}=5\sqrt{2}"

The side of the square is 5.


"slope_{AC}=\dfrac{4-(-1)}{3-(-2)}=1"

Then


"slope_{BD}=-1"

The equation of the "BD" is "y=-x+b."

The center of the square is "O(\dfrac{3-2}{2},\dfrac{4-1}{2} )."

Substitute


"\dfrac{3}{2}=-\dfrac{1}{2}+b=>b=2"

"BD:y=-x+2"


The angle between "AB" and "BD" is "45\degree." Then the equation of "AB" is "x=-2."


"y=-(-2)+2=4"

"B(-2, 4)"

The angle between "CD" and "BD" is "45\degree." Then the equation of "CD" is "x=3."


"y=-3+2=-1"

"D(3, -1)"